OpenStudy (anonymous):

How do you graph f(x)=(square root)16+x^2? Please help!

OpenStudy (anonymous):

the equation will simplify to f(x)=4+x

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

was this the equation \[\sqrt{16+x ^{2}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

 based one the eqution simplified by Sweetangel f(x)=4+x

OpenStudy (anonymous):

yeah if you square root whats inside you get 4+x

OpenStudy (anonymous):

 this is what the example looks like in my book